1. A bulb drawn 0.4 A current from a 3 V battery in 6.5 minutes. Calculate the

(a) number of charge that flows in the circuit.

(b) resistance of the bulb filament.

(c) energy dissipated in the bulb.


1
Expert's answer
2022-02-07T15:21:23-0500

Part (a)

"I=\frac{Q}{t}"


"0.4=\frac{Q}{60×6.5}"

"Q=156C"


Part (b)

"V=IR"

"R=\frac{3}{0.4}=7.5\Omega"


Part (c)

"E=QV"

"=156\u00d73"

"=468J"